myininaya (myininaya):

did we ever figure out a pattern or recursive formula for what was it: 2,4,6,8,16,... and if you don't know what i'm talking about it's cool no worries lol

OpenStudy (anonymous):

time every thing by 2

myininaya (myininaya):

2n 2*1=2 2*2=4 2*3=6 2*4=8 2*5=10 does not work for the 5th term

OpenStudy (anonymous):

i could never figure it out =/

OpenStudy (anonymous):

the 16 throws everything off

myininaya (myininaya):

or the 6 2^1=2 2^2=4 2^3=8 2^4=16

OpenStudy (anonymous):

i just came up with a cheap idea.....one sec <.<

OpenStudy (anonymous):

never mine.. lol

myininaya (myininaya):

nt shasha :)

myininaya (myininaya):

i will reward everyone

myininaya (myininaya):

i wonder if there is a program that can detect a pattern

OpenStudy (anonymous):

ii might have gotten an answer <.< gonna test it in a sec

myininaya (myininaya):

go superjoemath!

OpenStudy (anonymous):

i dont deserve any praise lol, i basically abused the fact that we dont know what the end of the sequence looks like....so i made up my own sequence and made a formula for it >.>

myininaya (myininaya):

ok it can't be that bad lets hear it

OpenStudy (anonymous):

hahaha it worked >.> heres the formula, this gives the nth term of the sequence, and for n = 1,...,5, it gives the numbers above <.<

OpenStudy (anonymous):

holy, how did you come to that

myininaya (myininaya):

\[f_n=f_{n-1}+f_{n-2}+2^{n-2}, for n\ge 3, f_1=2,f_2=4\] i think this works too but its stupid let me look at yours

OpenStudy (anonymous):

honestly...id rather not explain, its a pretty ridiculous process >.>

myininaya (myininaya):

nope mine doesn't work

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=-%2823%2F12%29%2B7%2F2*2^n-3%2F2*3^n%2B5%2F12*4^n-1%2F20*5^n%2C+n+%3D+1 switch the n = x to whatever term, to make sure it works.

myininaya (myininaya):

thats crazy lol

OpenStudy (anonymous):

yay,1000th medal

myininaya (myininaya):

i think the sequence that guy was working on was not meant for him

OpenStudy (anonymous):

there are infinitely many formulas i can make to fit those first five numbers, all depending on how i want the tail of the sequence to turn out.

OpenStudy (anonymous):

if you know how to change: \[F_{n+2} = F_{n+1} + F_n\] into: \[F(n) = \frac{1}{\sqrt{5}}[(\frac{1+\sqrt{5}}{2})^{n}-(\frac{1-\sqrt{5}}{2})^{n}]\] then the process i used is similar-ish

myininaya (myininaya):

you win joe you are are the more evolved :(

OpenStudy (anonymous):

just because i know more about one topic doesnt mean i win >.< im super super super weak when it comes to multi-variable calc. no idea how to do anything with it >.< thats why i dont answer calc questions lol

myininaya (myininaya):

i forgot alot of math i guess if you don't use it you lose it :(

OpenStudy (anonymous):

hell, even single variable calc, scares me to death!

myininaya (myininaya):

like linear algebra and differential equations

OpenStudy (anonymous):

are you teacher?

OpenStudy (anonymous):

Linear Algebra is beast <.< the above technique is based off linear algebra

myininaya (myininaya):

i teach but don't tell no one

myininaya (myininaya):

joe what do you do

OpenStudy (anonymous):

im a lowly under grad >.< still working on my degree.

myininaya (myininaya):

lol

OpenStudy (anonymous):

nothing but a college dropout trying second time <.<

myininaya (myininaya):

somehow i got my masters

OpenStudy (anonymous):

master in math?

myininaya (myininaya):

yes

OpenStudy (anonymous):

thats awesome O.o

OpenStudy (anonymous):

im jealous o.O

myininaya (myininaya):

but i don't think im a math person i'm more interested in cryptography

OpenStudy (anonymous):

like john nash?

myininaya (myininaya):

i want to get good at number theory and group theory and linear algebra i only had elementary number theory and i can't remember linear algebra and group theory much i really wish my school focused more on number theory that would had been amazinf

OpenStudy (anonymous):

im thinking about taking a class in cryptography, i hope i can fit it in to my degree plan.

OpenStudy (anonymous):

my school has a horrible math dept >.< i have to buy books and read to keep up with the rest of the world >.>

myininaya (myininaya):

i should had did that

myininaya (myininaya):

i think they can only teach you so much you don't come out learning everything you should

OpenStudy (anonymous):

85-90% of students in the math dept are education majors, they only care about what they will need to teach in high school and thats it

myininaya (myininaya):

i don't want to teach i think it would be cool to work at the nsa

OpenStudy (anonymous):

linear is last math class I am going to take

OpenStudy (anonymous):

Linear is amazing >.>

myininaya (myininaya):

joe is really good at linear ask him all the questions you like

myininaya (myininaya):

i have food peace

OpenStudy (anonymous):

next fall I will take diffiq and in spring Linear algebra

OpenStudy (anonymous):

im taking Number Theory, Matrix Theory, and Diff Eq next semester...im a little worried.